OpenStudy (anonymous):

I have to find the most general antiderivative f(X)= pi^6

OpenStudy (anonymous):

As you know, \(\pi^6\) is a constant. Therefore, the variable x in the anti-derivative has teh power of 1. Have you figure it out from here?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Can you share the answer?

OpenStudy (anonymous):

My answer is \(\pi^6x+C\)

OpenStudy (anonymous):

i had pi^6x

OpenStudy (anonymous):

so most general antiderivative is just another way to say "find the derivative"?

OpenStudy (anonymous):

Well done.

OpenStudy (anonymous):

i meant the antiderivative

OpenStudy (anonymous):

Yeah, almost.

OpenStudy (anonymous):

For example, your answer for this problem is \(\pi^6x\). It is correct. However, the most general antiderivative will be \(\pi^6x+c\). Because \(\int(\pi^6)dx =\pi^6x\). But \(\int(\pi^6)dx =\pi^6x+1\) or \(\pi^6x+2\) and so on, right?

OpenStudy (anonymous):

So they use \(c\) to write the most general antiderivative, always, after find the antiderivative

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

There are soee more examples here. http://www-math.mit.edu/~djk/calculus_beginners/chapter16/section01.html

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are welcome